OpenStudy (anonymous):

Two pipes together can fill a reservoir in 6 hours and 40 minutes. Find the time each alone will take to fill the reservoir if one of the pipes can fill it in 3 hours less than the other.

hero (hero):

\[\frac{AB}{A + B} = t\]

hero (hero):

A = Reservoir 1 B = Reservoir 2 A = x B = x - 3 \[\frac{x(x-3)}{x + (x-3)} =6\frac{2}{3}\]

hero (hero):

Solve for x

OpenStudy (anonymous):

How did you get 6 ?

hero (hero):

It was given in the problem. We still have to find x

OpenStudy (anonymous):

okay..

hero (hero):

Hang on a sec, I have to re-compute that last part

hero (hero):

x = 15 x - 3 = 12

hero (hero):

That's what you should get

OpenStudy (anonymous):

Okay , I'm trying to comprehend , lol

OpenStudy (anonymous):

Where did 15 come from ? how did we solve for x

hero (hero):

It's a long story

OpenStudy (anonymous):

Hmmm.